A 60 kg person is riding a roller coaster with radius of 6 m At top of curve the normal force acting on the person is 120N Determine the speed of the car at this point
Answer #1

Normal force = m v^2 / r

F = m v^2 / r

v = sqrt( F r / m) = sqrt( 120 * 6 / 60 ) = 3.46 m/s

=====================================================

Speed of the car at this point = 3.46 m/s